# End-to-End Dexterous Arm-Hand VLA Policies via Shared Autonomy: VR Teleoperation Augmented by Autonomous Hand VLA Policy for Efficient Data Collection

### v1

@@ 3.2.2 Tactile-based DexGrasp-VLA $\pi_{\text{hand}}$
Although the LSTM policy provides robust low-level force adaptation, it lacks visual perception and scene understanding. To incorporate visual context and enable task-aware grasping (e.g., timing to grasp the target object in clutter), we use the data autonomously collected by the LSTM policy to train a multimodal VLA policy that integrates tactile sensing for robust dexterous grasping.
@@
Tactile Feature Extraction
Raw tactile readings from the dexterous hand sensors are captured as a high-dimensional tensor $F_{\text{raw}}\in\mathbb{R}^{10\times 12\times 3}$ per fingertip, representing force values along the $x$, $y$, and $z$ axes of a spatial grid over the surface of the fingertip. Directly feeding such high-dimensional data into a VLA policy is computationally inefficient. To overcome this, we derive two complementary tactile features that provide a compact yet rich representation of contact conditions.
@@
The encoder comprises three convolutional layers. with $3\times 3$ kernels and a stride of 2, using 32, 64, and 128 filters, respectively. Each layer is followed by batch normalization and a ReLU activation, progressively reducing the spatial resolution while increasing feature depth. This hierarchical encoding results in a $2\times 2\times 128$ feature map, which is flattened and projected to a 128-dimensional latent vector per fingertip. The decoder mirrors this architecture using transposed convolutions to reconstruct the original tactile image from the latent code.
@@ 3.3.1 VR-Based Arm Teleoperation System
$\min_{\dot{q}}\sum_{i=1}^{N}w_{i}\left\|J_{i}(q)\dot{q}+e_{i}(q)\right\|^{2}.$ (7)
$\text{s.t.}\quad l\leq C(q)\dot{q}\leq u$
@@
where $\Delta t$ is the sampling time of the control loop, and $q_{t}$ is the previous joint angles. This process iterates until the task errors $\|e_{i}(q)\|$ converge below a threshold or a terminal condition is met. The VR-based teleoperation system captures the motions of the controller, mapping them in real time to the robotic arm’s motion. This setup enables natural and responsive arm control in 3D space, significantly improving demonstration efficiency and precision.
@@ 3.3.2 Coordinated Arm-Hand Data Collection
For data collection, we engineered a multi-thread control architecture that enables seamless integration of human teleoperation (90 Hz) with autonomous policy execution (30 Hz), as well as streaming and saving time-synchronized data (30 Hz). In this setup, the human operator controls the arm’s 6-DoF end-effector motion through a VR interface while the DexGrasp-VLA policy runs concurrently, generating appropriate grasping actions based on real-time visual and tactile feedback. This parallel execution mechanism preserves natural temporal coordination between arm’s reaching and hand’s grasping phases, enabling efficient collection of synegized arm-hand demonstrations.
The collected datasets contain temporally synchronized observations and actions from both control sources, formally represented as:
$\begin{gathered}\mathcal{D}_{\text{uni}}=\{(o_{t}^{\text{uni}},a_{t}^{\text{arm}},a_{t}^{\text{hand}})\}_{t=1}^{T},\\ a_{t}^{\text{arm}}\sim p_{\text{teleop}},\quad a_{t}^{\text{hand}}\sim\pi_{\text{hand}}(\cdot\mid o_{t}^{\text{hand}}),\end{gathered}$ (9)
@@ 3.4 Learning End-to-End Arm-Hand VLA Policy $\pi_{\text{uni}}$
Building upon the arm-hand demonstration data collected through shared autonomy, we perform SFT of a pre-trained VLA model for learning an arm-hand coordinated dexterous grasping policy $\pi_{\text{uni}}(\mathcal{A}_{t}^{\text{uni}}\mid o_{t}^{\text{uni}})$. A key challenge lies in effectively handling the distinct characteristics of arm and hand motion. Our prior research efforts found that a vanilla version of SFT would only lead to high performance, because of which we invented the technical improvement that has no structural alteration as for the fine-tuning stage – the Arm-Hand Feature Enhancement module that extends the base architecture with explicit mechanisms to capture both shared task context and unique individual features for arm and hand dynamics, respectively.
@@ 3.4.1 Arm-Hand Feature Enhancement
Reasoning from the first principle, we developed the Arm-Hand Feature Enhancement module to resolve this, which explicitly models the complementary roles of macro-movement (arm) and micro-manipulation (hand). The shared representation $\bm{z_{t}^{\text{share}}}$ is passed through two dedicated multi-layer perceptrons (MLPs): $\bm{E_{\text{arm}}}$ and $\bm{E_{\text{hand}}}$. These produce limb-specific latent features $\bm{z_{t}^{\text{arm}}}\in\mathbb{R}^{d_{a}}$ and $\bm{z_{t}^{\text{hand}}}\in\mathbb{R}^{d_{h}}$, respectively.
@@ 3.4.2 Learning Objective
where $\mathcal{A}_{t}^{\tau}=\tau\mathcal{A}_{t}+(1-\tau)\epsilon$ is a noise-corrupted action chunk and $u(\mathcal{A}_{t}^{\tau}\mid\mathcal{A}_{t})=\epsilon-\mathcal{A}_{t}$ is the target vector field.
@@
To reinforce the disentanglement and specialization of arm and hand features, we introduce two auxiliary losses:
@@ 4.1.1 Manipulation Platform
Our experimental platform consists of an integrated robotic manipulator with a hand and a multi-sensor perception system, as shown in Fig. 8, which provides the necessary physical and sensory capabilities to develop and evaluate our shared autonomy and learning algorithms for dexterous manipulation.
@@ 4.1.2 Datasets
Figure 11: Grasping and placing various objects via end-to-end (arm-hand) VLA policies with a dexterous hand. (See Appendix 8 for more real robot tasks.)
@@ 4.2 Main Results
4.2.1 Grasping Performance of the VLA Policy $\pi_{\text{hand}}$
@@ 4.3 Ablation study
Figure 12: Effectiveness of tactile-based DexGrasp-VLA $\pi_{\text{hand}}$ and robust grasping performance with tactile sensing under camera occlusion: (a) Without the sense of touch (tactile), the object slips while camera is occluded; (b) With dual tactile features, the policy sustains firm and stable grasp throughout the test without any slippage, even under perturbations such as bumping against the table (9s) and intense up-down shaking (9s-12s).
@@ 4.3.2 Effectiveness of Arm-Hand Feature Enhancement in the VLA Policy $\pi_{\text{uni}}$
We conducted a comprehensive ablation study in 3 experimental settings (Table 3) to evaluate the effectiveness of the proposed Arm-Hand Feature Enhancement module. The enhanced policy $\pi_{\text{uni-enhance}}$ is compared against the baseline $\pi_{\text{uni-origin}}$(initialized from $\pi_{0}$ [5]), with both models fine-tuned on the same dataset $\mathcal{D}_{\text{uni}}$. Results across three settings highlight consistent performance gains, and the comparison result is shown in Fig. 14. On the main Xhand platform, $\pi_{\text{uni-enhance}}$ achieves a 95% success rate compared to 88% for the baseline, confirming improved coordination and grasp stability. When transferred to the RY-H2 dexterous hand [1], the enhanced policy outperforms the baseline by 10% (81% vs. 71%), showing that the framework is applicable to different hardware that varies in kinematics and actuation (See more specifications of dexterous hands in Appendix 7.2).
@@ 5.1 Discussion
In this work, with limited time and resources, our evaluation focuses on grasping tasks – a balanced choice since grasping serves as a fundamental and ubiquitous primitive in robotic manipulation, providing a well-defined yet challenging benchmark to validate shared autonomy, multimodal policy learning, and corrective teleoperation under diverse physical contact conditions. Despite these promising results as a zero-to-one proof of concept, limitations remain: the current system has not been tested in more complex long-horizon manipulation, such as tool use or in-hand reorientation; the full tactile integration, while critical for grasp stability, remains challenging due to noise and feature misalignment with arm movements (reaching without touching); and corrective mechanism and continual improvements still rely on human intervention, which may limit the scalability and fully autonomous approaches are deemed more promising at a larger scale.
@@ 5.2 Future Work
The proposed architecture and methodology are designed in a modular manner for possible extensions. The shared autonomy interface naturally separates high-level human guidance from low-level execution, allowing new VLA copilots to be modularly integrated. The policy structure, particularly the arm–hand feature enhancement, is task-agnostic and can adapt to diverse demonstrations, while the corrective teleoperation framework applies readily to new tasks.
Future work would need a community effort to improve the VLA paradigm in multiple directions. This framework can be extended to a broader set of manipulation skills, including object re-orientation, insertion, and long-horizon tasks, by developing or combining additional specialized AI-driven copilots. We will also investigate advanced tactile representation learning techniques to better fuse tactile, visual, and proprioceptive inputs, for example, through selective routing or conditional gating, and explore auxiliary prediction losses inspired by the feature enhancement module.
Furthermore, future work can focus on autonomous error detection and composite recovery strategies. First, the human-in-the-loop teleoperation correction phase can be replaced by an autonomous error recovery module. Reinforcement learning and self-supervised adaptation can be used to train a model to autonomously detect and recover from errors, thus reducing dependence on human intervention. Second, the coordination and switching of control authority between the human and the robot can be dynamically and seamlessly adjusted based on the perception of the context, such as task difficulty, environmental safety, and human state. This perceptual capability would be achieved via VLMs, and allow the system to autonomously determine when it can operate autonomously error recovery by itself, and when it needs human input or intervention.
@@ 5.3 Conclusion
Moreover, it also suggests a potential new step change in VLA-based physical AI systems, which can be underpinned by VLMs’ perceptual understanding ability to match human-level cognition and thus to deal with complicated scenarios, where a repertoire of autonomous robot skills and human-in-the-loop mechanism can be synergistically combined for deploying such technologies in real-world industrial sectors.
@@ 6 Contributions & Acknowledgements
We sincerely thank Jianran Liu, Yufei Liu, Lisuo Li for their generous and kind support on hardware setup, and Wangli Peng for his support in providing hardware accessories. We thank Heyang Xu and Ke Fan for their help on the early prototyping of the VLA systems, and Jieyi Zhang for his early development of the LSTM-based grasping policies.
We extend our sincere gratitude to Wenjia Zhu, for fostering our focus on high-value research. We are very thankful to Yonghui Wu for his strategic direction in shaping research priorities. Furthermore, special acknowledgment goes to the Department Head, Hang Li, for his enabling and organizational support throughout this project.
@@ 7.2 Hardware Platforms and Real-World Deployment
Figure 16: Two different robotic hands used in the study for validating the Shared Autonomy framework for training end-to-end VLA models: (a) XHAND1 hand [2], and (b) RY-H2 hand [1].
This research uses two types of dexterous robotic hands: (1) Xhand, a high-performance 12-DoF dexterous robotic hand, as the main hardware; and (2) RY-H2, a fast 11-DoF dexterous hand (6 active DoF and 5 under-actuated DoF) with a quick open-close cycle of 0.4s.
@@
During the evaluation phase on the physical robot system, the following configurations are applied:
@@ 8.1 Effectiveness of Shared Autonomy Data Collection
Full Teleoperation 90/hour/person 80/hour/person - -
@@ 8.2.1 Effectiveness of Tactile Sensing in $\pi_{\text{uni}}$
This performance degradation is likely due to the different functional roles for controlling the arm and the hand respectively. Unfortunately, due to the lifetime and closure of this project, no continued efforts could be pursued for further investigation. To the best of our knowledge, our current understanding is shared as follows. The arm primarily executes reaching motions that rely more on visual and proprioceptive feedback for spatial guidance, while tactile signals are most relevant for fine-grained grasping and in-hand manipulation. Uniformly incorporating tactile input throughout the entire arm-hand trajectory may introduce irrelevant information like “noises” during arm movement phases, particularly from incidental environmental contacts (e.g., table collisions or unintended fingertip brushing) that occur during reaching. These transient and often misleading tactile signals appear to interfere with the policy’s ability to maintain robust arm-centric coordination.

### v2

@@ 3.2.2 Tactile-based DexGrasp-VLA $\pi_{\text{hand}}$
Although the LSTM policy provides robust low-level force adaptation, it lacks visual perception and scene understanding. To incorporate visual context and enable task-aware grasping (e.g., timing to grasp the target object in clutter), we use the data autonomously collected by the LSTM policy to train a multimodal VLA policy that integrates tactile sensing for robust dexterous grasping, as detailed in Fig. 4.
Tactile Feature Extraction
@@
The raw tactile readings from the dexterous hand sensors are captured as a high-dimensional tensor $F_{\text{raw}}\in\mathbb{R}^{10\times 12\times 3}$ per fingertip, representing force values along the $x$, $y$, and $z$ axes of a spatial grid over the surface of the fingertip. Directly feeding such high-dimensional data into a VLA policy is computationally inefficient. To overcome this, we derive two complementary tactile features that provide a compact yet rich representation of contact conditions, as shown in Fig. 4.
@@
The encoder comprises three convolutional layers: $3\times 3$ kernels and a stride of 2, using 32, 64, and 128 filters, respectively. Each layer is followed by batch normalization and a ReLU activation, progressively reducing the spatial resolution while increasing feature depth. This hierarchical encoding results in a $2\times 2\times 128$ feature map, which is flattened and projected to a 128-dimensional latent vector per fingertip. The decoder mirrors this architecture using transposed convolutions to reconstruct the original tactile image from the latent code.
@@ 3.3.1 VR-Based Arm Teleoperation System
$\min_{\dot{q}}\sum_{i=1}^{N}w_{i}\left\|J_{i}(q)\dot{q}+e_{i}(q)\right\|^{2},\hskip 5.69054pt\text{s.t.}\quad l\leq C(q)\dot{q}\leq u.$
@@
where $\Delta t$ is the sampling time, and $q_{t}$ are the previous joint angles. This process iterates until the task errors $\|e_{i}(q)\|$ are below a threshold or a terminal condition is met. The VR-based teleoperation system captures and maps the motions of the VR controller in real time to the robotic arm’s motion. This setup enables natural and responsive arm control in 3D space, significantly improving demonstration efficiency and precision.
@@ 3.3.2 Coordinated Arm-Hand Data Collection
For data collection, we engineered a multi-thread control architecture that enables seamless integration of human teleoperation (90 Hz) with autonomous policy execution (30 Hz), as well as streaming and saving time-synchronized data (30 Hz). In this setup, the human operator controls the arm’s 6-DoF end-effector motion through a VR interface while the DexGrasp-VLA policy runs concurrently, generating appropriate grasping actions based on real-time visual and tactile feedback. This parallel execution mechanism preserves natural temporal coordination between arm’s reaching and hand’s grasping phases, enabling efficient collection of synegized arm-hand demonstrations. The collected datasets contain temporally synchronized observations and actions from both control sources, formally represented as:
$\begin{gathered}\mathcal{D}_{\text{uni}}=\{(o_{t}^{\text{uni}},a_{t}^{\text{arm}},a_{t}^{\text{hand}})\}_{t=1}^{T};\hskip 5.69054pta_{t}^{\text{arm}}\sim p_{\text{teleop}},\quad a_{t}^{\text{hand}}\sim\pi_{\text{hand}}(\cdot\mid o_{t}^{\text{hand}}),\end{gathered}$ (8)
@@ 3.4 Learning End-to-End Arm-Hand VLA Policy $\pi_{\text{uni}}$
Building upon the arm-hand demonstration data collected through shared autonomy, we perform SFT of a pre-trained VLA model to learn an arm-hand coordinated dexterous grasping policy $\pi_{\text{uni}}(\mathcal{A}_{t}^{\text{uni}}\mid o_{t}^{\text{uni}})$. A key challenge lies in effectively handling the different and complementary characteristics of arm and hand motion. Our prior research efforts found that a vanilla version of SFT would only lead to limited performance, primarily because it fails to explicitly disentangle and represent the unique dynamics of the arm and hand, while also capturing their shared task context.
To overcome this limitation, we developed the technical improvement that has no structural alteration as for the fine-tuning stage – the Arm-Hand Feature Enhancement module. It extends the base architecture with explicit mechanisms to capture both shared contextual information relevant to coordinated arm-hand tasks, and distinct individual features that characterize arm and hand dynamics, respectively. Hence, it enables more effective and robust arm-hand coordination during dexterous grasping, for example, even under visual occlusion of one camera.
@@ 3.4.1 Arm-Hand Feature Enhancement
Reasoning from the fundamental differences in kinematic and dynamic interactions and functional roles, we developed the Arm-Hand Feature Enhancement module to resolve this issue, which explicitly models the complementary roles of macro-movement (arm) and micro-manipulation (hand). The shared representation $\bm{z_{t}^{\text{share}}}$ is passed through two dedicated multi-layer perceptrons (MLPs): $\bm{E_{\text{arm}}}$ and $\bm{E_{\text{hand}}}$. These produce limb-specific latent features $\bm{z_{t}^{\text{arm}}}\in\mathbb{R}^{d_{a}}$ and $\bm{z_{t}^{\text{hand}}}\in\mathbb{R}^{d_{h}}$, respectively.
@@ 3.4.2 Learning Objective
where $\mathcal{A}_{t}^{\tau}=\tau\mathcal{A}_{t}+(1-\tau)\epsilon$ is a noise-corrupted action chunk, $u(\mathcal{A}_{t}^{\tau}\mid\mathcal{A}_{t})=\epsilon-\mathcal{A}_{t}$ is the target vector field.
@@
Two auxiliary losses are used to reinforce the disentanglement and specialization of arm and hand:
@@ 4.1.1 Manipulation Platform
Our experimental platform, consisting of an integrated robotic manipulator with a hand and a multi-sensor perception system (as shown in Fig. 8), provides the necessary physical and sensory capabilities to develop and evaluate our shared autonomy and learning algorithms for dexterous manipulation.
@@ 4.2 Main Results
4.2.1 Grasping Performance of the DexGrasp-VLA Policy $\pi_{\text{hand}}$
@@ 4.2.2 Grasping Performance of the VLA Policy $\pi_{\text{uni}}$
Figure 11: Grasping and placing various objects via end-to-end (arm-hand) VLA policies with a dexterous hand. (See Appendix 8 for additional real robot tests.)
@@ 4.3 Ablation study
Figure 13: Effectiveness of tactile-based DexGrasp-VLA $\pi_{\text{hand}}$ and robust grasping performance with tactile sensing under camera occlusion: (a) Without the sense of touch (tactile), the object slips while camera is occluded; (b) With dual tactile features, the policy demonstrates a sustained, stable grasp throughout the test without any slippage, even under external perturbations (bumping against the table at 9s, and intense up-down shaking from 9s-12s).
@@ 4.3.2 Effectiveness of Arm-Hand Feature Enhancement in the VLA Policy $\pi_{\text{uni}}$
We conducted a comprehensive ablation study in 3 experimental settings, as shown in Table 3, to evaluate the effectiveness of the proposed Arm-Hand Feature Enhancement module. The enhanced policy $\pi_{\text{uni-enhance}}$ is compared against the baseline $\pi_{\text{uni-origin}}$(initialized from $\pi_{0}$ [5]), with both models fine-tuned on the same dataset $\mathcal{D}_{\text{uni}}$. Results across three settings highlight consistent performance gains, and the comparison result is shown in Fig. 15. On the main Xhand platform, $\pi_{\text{uni-enhance}}$ achieves a 95% success rate compared to 88% for the baseline, confirming improved coordination and grasp stability. When transferred to the RY-H2 dexterous hand [1], the enhanced policy outperforms the baseline by 10% (81% vs. 71%), showing that the framework is applicable to different hardware that varies in kinematics and actuation (See more specifications of dexterous hands in Appendix 7.2).
@@ 5.1 Discussion
In this work, we focus on grasping tasks as a fundamental capability in robotic manipulation. The high degree-of-freedom and complex multi-contact dynamics of five-finger dexterous hands impose a significant challenge, where numerous factors, from tactile sensing to arm-hand coordination, are deeply intertwined. Therefore, grasping serves as an ideal testbed, providing a well-defined yet challenging benchmark to validate shared autonomy, multimodal policy learning, and corrective teleoperation under diverse physical contact conditions. It allows for a methodological focus to systematically isolate and investigate the role of each core component, such as a more structured algorithmic integration of multimodal sensorimotor pathways for tactile sensing and different techniques for shared autonomy, without involving the extra complexity of dexterous in-hand manipulation. This focused approach enables a clear validation of our framework, establishing a foundational understanding of multimodal policy learning that paves the way for developing and scaling up to more sophisticated large-scale foundational models with human-level dexterity.
Despite these promising results as a zero-to-one proof of concept, limitations remain: the current system has not been tested in more complex long-horizon manipulation, such as tool use or in-hand reorientation; the full tactile integration, while critical for grasp stability, remains challenging due to noise and feature misalignment with arm movements (reaching motions without touching); and corrective mechanism and continual improvements still rely on human intervention, which limits the scalability and thus fully autonomous approaches are clearly more promising at a larger scale.
@@ 5.2 Future Work
The proposed architecture and methodology are designed with inherent extensibility, which allows a cohesive way for future upgrade with more advanced hardware, e.g., the VLA copilot system and the teleoperation are combinatorial. The shared autonomy interface naturally separates high-level human guidance from low-level execution, allowing new VLA copilots for various skills to be integrated within the same framework, for example, more sophisticated dexterous in-hand manipulation with multi-finger hands. Similarly, the policy structure, particularly the arm–hand feature enhancement, is task-agnostic and can adapt to diverse demonstrations, while the corrective teleoperation framework applies readily to new tasks.
Future work would call for a community effort to improve the VLA paradigm in multiple directions. This framework can be extended to a broader set of manipulation skills, including object re-orientation, insertion, and long-horizon tasks, by developing or combining additional specialized AI-driven copilots. We will also investigate advanced tactile representation learning techniques to better fuse tactile, visual, and proprioceptive inputs, for example, through selective routing or conditional gating, and explore auxiliary prediction losses inspired by the feature enhancement module.
Furthermore, future work can focus on autonomous error detection and composite recovery strategies. First, the human-in-the-loop teleoperation correction phase can be replaced by an autonomous error recovery module. Reinforcement learning and self-supervised adaptation can be used to train a model to autonomously detect and recover from errors, thus reducing the dependence on human intervention. Second, the coordination and switching of control authority between the human and the robot can be dynamically and seamlessly adjusted based on the perception of the context, such as task difficulty, environmental safety, and human state. This perceptual capability would be achieved via VLMs, and allow the system to autonomously determine when it can autonomously operate error recovery by itself, and when it needs human input or intervention.
@@ 5.3 Conclusion
Moreover, it also suggests a potential new step change in VLA-based physical AI systems, which can be leveraged by VLMs’ perceptual understanding ability to match human-level cognition, and thus to deal with complicated scenarios, where a repertoire of autonomous robot skills and human-in-the-loop mechanism can be synergistically combined for deploying such technologies in real-world industrial sectors.
@@ 6 Contributions & Acknowledgements
We sincerely thank Jianran Liu, Yunfei Liu, Lisuo Li for their generous and kind support on hardware setup, and Wanli Peng for his support in providing hardware accessories. We thank Heyang Xu and Ke Fan for their help on the early prototyping of the VLA systems, and Jieyi Zhang for his early development of the LSTM-based grasping policies.
We extend our sincere gratitude to Wenjia Zhu, for encouraging the focus on high-value research. We are very thankful to Yonghui Wu for his strategic direction in shaping research priorities. Furthermore, our special acknowledgment goes to the Department Head, Hang Li, for his substantial enabling and organizational support throughout this project.
@@ 7.2 Hardware Platforms and Real-World Deployment
Figure 17: Cross-embodiment validation using two different robotic hands for both the Shared Autonomy framework and the training of end-to-end VLA models: (a) XHAND1 hand [2], and (b) RY-H2 hand [1].
As the research community embraces a range of different designs of hardware, we validated the applicability of our proposed framework using two different dexterous robotic hands which are quite representative (one is fully actuated and the other is underactuated): (1) Xhand, a high-performance 12-DoF dexterous robotic hand with fingertip tactile sensing, as the main hardware; (2) RY-H2, a fast 11-DoF dexterous hand (6 active DoF and 5 under-actuated DoF) with joint current sensing featured by a quick open-close cycle of 0.4s.
@@
During the evaluation phase on these physical robot hardware, the same setting was configured:
@@ 8.1 Effectiveness of Shared Autonomy Data Collection
Full Teleoperation 90/hour/person 80/hour/person NA NA
@@ 8.2.1 Effectiveness of Tactile Sensing in $\pi_{\text{uni}}$
This performance degradation is likely due to the different functional roles for controlling the arm and the hand respectively. To the best of our knowledge, we hypothesize that the arm primarily executes reaching motions that rely more on visual and proprioceptive feedback for spatial motions, while tactile signals are most relevant for fine-grained grasping and in-hand manipulation. Uniformly incorporating tactile input throughout the entire arm-hand trajectory may introduce irrelevant information like “noises” during arm movement phases, particularly from incidental environmental contacts (e.g., table collisions or unintended fingertip brushing) that occur during reaching. These transient and often misleading tactile signals appear to interfere with the policy’s ability to maintain robust arm-centric coordination.